Graph the system of equations y=x^4-3x³ and y=2x-6. Find the points of intersection. Give an exact answer. PLEASE HELP ME! I don't know what to do!!
Answer by HyperBrain(694)

y=x^4-3x³ and y=2x-6
Replace the y in the first equation by 2x-6
2x-6=x4-3x3
x4-3x3-2x+6=0
(x4-3x3)-(2x-6)=0
x3(x-3)-2(x-3)=0
(x3-2)(x-3)
Let x-3=0, then x=3
Let x3-2=0, then x=cube root of 2>>get calculator... i lost mine
Now, we have the x-coordinates of the two intersections. To find the corresponding y-coordinates, use y=2x-6.
So,
if x=3, y=2(3)-6=6-6=0
if x=2*(cube root of 2)-6>>please! calculator again....
So the points of intersections are (3,0) and (cube root of 2, 2*(cube root of 2)-6)
Power up,
HyperBrain!
Stop this mess! http://uselessplastic.wordpress.com/
http://uselessplastic.multiply.com/